OpenStudy (anonymous):

Write an indirect proof proving only one angle of an obtuse triangle is greater than 90°.

OpenStudy (anonymous):

Indirect proof is a proof of contradiction. List the things you know abot an obtuse triangle - It's a triangle so must have interior angles that total 180º \[ a+b+c=180º\] - An obtuse must have one angle that is larger than 90º and less than 180º \[90º<a<180º \ \ \ \ \mathrm{or} \ \ \ \ 90º<b<180º \ \ \ \ \mathrm{or} \ \ \ \ 90º<c<180º\] Now you assume that at least 2 angles of the triangle have angles greater than 90º Can you write an equation or expression for that part?

OpenStudy (anonymous):

I don't think I can.

OpenStudy (anonymous):

well put somethin' down and we'll see where it takes us ^_^

OpenStudy (anonymous):

I BELIEVE IN YOU

OpenStudy (anonymous):

where do you get the degrees sign

OpenStudy (anonymous):

are you on a mac or a pc?

OpenStudy (anonymous):

pc

OpenStudy (anonymous):

alt+0176 on the number pad

OpenStudy (anonymous):

90 degrees> a> 180 degrees and 90 degrees> b> 180 something like that?

OpenStudy (anonymous):

exactly! (well, almost. it's 90º<a<180º and 90º<b<180º - the inequalities open to the right to make it say "90 degrees is less than a, a is less than 180 degrees," but I knew what you meant. The AND was the important part) So if a AND b are both greater than 90º, what happens to c when we try to sum up the angles \[(a+b)+c=180\]

OpenStudy (anonymous):

c will make the total sum of the degrees over 180.

OpenStudy (anonymous):

Exactly!! If c has any non-negative value then the left hand side of the equation will be greater than 180º (also, alt+0176 means you hold down the alt key then press 0176 on the number pad, then release the alt key)

OpenStudy (anonymous):

You can show that graphically by 

OpenStudy (anonymous):

I understand that, thank you. But ordering the facts into a logical proof is also giving me a hard time.

OpenStudy (anonymous):

^_^ welcome You start with the things that must be true for every triangle  \[\textrm{For every triangle} \\ \\ \ \\ \ a+b +c\overset{!}{=}180º\] \[ \textrm{For every obtuse triangle} \\ \ \\ \ 90º<a<180º\] \[\textrm{Assumption: both a and b are greater than 90º} \\ \ \\ \ 90º<a<180º \ \ \ ; \ \ \ 90º<b<180º\] \[\textrm{If true} \\ \ \\ \ 180º<a+b<360º\] \[\textrm{Then for} \\ \ \\ \ (a+b)+c\overset{!}{=}180º \\ \ \\ \ \\ c<0\] \[ \textrm{This makes no physical sense. Thus only} \\ \ \\ \ \\ 90º<a<180º \\ \ \\ \ \textrm{can be true. There exists only one obtuse angle in an obtuse triangle. QED.}\] the equals sign means "must be equal to." You could also use the triple equals sign \[ \equiv\]

OpenStudy (anonymous):

I think I know what to do now. Thanks for your help. :)

OpenStudy (anonymous):

welcome ^_^ Good luck with the writeup!